Find a symbolic representation for f^-1(x) show work?
F(x)=2x^2-3

Answer by rothauserc(4718)

we are looking for the inverse of the function
F(x) = 2x^2 -3
rewrite with F(x) as y
y = 2x^2 -3
solve for x
add 3 to both sides of =
y +3 = 2x^2
divide both sides of = by 2
x^2 = (y+3) / 2
x = square root( (y+3) / 2 )
now interchange the x's and y's
y = square root( (x+3) / 2 )
therefore
f^-1(x) = square root( (x+3) / 2 )